A box contains 20 chips. 8 of these chips have a value of $5 and the others have no value. We randomly pick 3 chips from the box and putting them back after each draw.

A = Number of chips of $5 after 2 draws.

What is the distribution of A?
Answer #1

Solution

Back-up Theory

Number of ways of selecting r things out of n things is given by nCr = (n!)/{(r!)(n - r)!}….....................................…(1)

Values of nCr can be directly obtained using Excel Function: Math & Trig COMBIN….......................................…. (1a)

Probability of an event E, denoted by P(E) = n/N ……......................................................................…………………(2)

where n = n(E) = Number of outcomes/cases/possibilities favourable to the event E and

N = n(S) = Total number all possible outcomes/cases/possibilities.

Now, to work out the solution,

Vide (1),

Number of ways of selecting 3 chips out of 20 chips = 20C3 = 1140. [vide (1a)]

So, vide (2), N = 1140 ............................................................................................................................................. (3)

For a single draw of picking 3 chips, let A1 = Number of chips of $5 in the first draw.

A1 = 0 => all 3 chips are $0 chips => Number of ways of selecting 3 chips out of 12 $0 chips = 12C3 = 220. [vide (1a)].

This is n vide (1). Hence , vide (1) and (3), P(A1 = 0) = 220/1140 = 0.1930

A1 = 1 => 1 chips is $5 and 2 are $0 chips => Number of ways of selecting 1 chip out of 8 $5 chips and 2 chips out of 12 $0 chips = (8C1) x (12C2)= 8 x 66. [vide (1a)] = 528. This is n vide (1). Hence , vide (1) and (3), P(A1 = 1) = 528/1140 = 0.4632

A1 = 2 => 2 chips are $5 and 1 is $0 chips => Number of ways of selecting 2 chips out of 8 $5 chips and 1 chip out of 12 $0 chips = (8C2) x (12C1)= 28 x 12. [vide (1a)] = 336. This is n vide (1). Hence , vide (1) and (3), P(A1 = 2) = 336/1140 = 0.2947

A1 = 3 => all 3 chips are $5 chips => Number of ways of selecting 3 chips out of 8 $5 chips = 8C3 = 56. [vide (1a)]

This is n vide (1). Hence , vide (1) and (3), P(A1 = 3) = 56/1140 = 0.0492

If A2 = Number of chips of $5 in the second draw, probability distribution of A2 is identical to that of A1 as given above.

Now, A = A1 + A2 and P(A) = P(A1) x P(A2).

All possible values of A and the corresponding probabilities are given in the following Table.

A1

A2

A

P(A)

0

0

0

0.0372

0

1

1

0.0894

0

2

2

0.0569

0

3

3

0.0095

1

0

1

0.0894

1

1

2

0.2145

1

2

3

0.1365

1

3

4

0.0228

2

0

2

0.0569

2

1

3

0.1365

2

2

4

0.0869

2

3

5

0.0145

3

0

3

0.0095

3

1

4

0.0228

3

2

5

0.0145

3

3

6

0.0024

From the above we derive the probability distribution of A as given below;

A

P(A)

0

0.0372

1

0.1788

2

0.3283

3

0.2920

4

0.1324

5

0.0290

6

0.0024

Total

1.0000

Answer

DONE